# Shortest-path percolation on scale-free networks

### v1

@@ IV.2 Shortest-path percolation process homogenizes the network drastically
where $\left\langle\cdot\right\rangle$ denotes the ensemble average across different realizations, $\phi_{o}$ is the asymptotic limit of the ratio, $\delta$ is the scaling exponent, and $a$ is a constant. Eq. (14) implies that if a network still holds its scale-freeness at $r=r^{\dagger}$, we expect $k_{\max}^{\dagger}$ will scale with the size of the network $N$ in the same way as $k_{\max}$ does. In this case, we expect $\phi_{o}$ is non-zero. On the other hand, we expect to obtain $\phi_{o}=0$ if the SPP process drastically homogenize the network.
We sample the observables at $r=r^{\dagger}$ using two methods: we collect the observables (i) at a fixed value of $r^{\dagger}=q$ for each realization or (ii) at $r^{\dagger}=qr_{c}^{*}(N)$ for each realization, where $q\in[0,1]$. Here, $q$ is a tunable parameter to decide on how far from the transition point we would like to sample the observables. We choose q=0.7, 0.8,q=0.7,\ 0.8, and $0.9$ to ensure that we are sampling the observables at the supercritical regime. For convenience, we add a subscript to differentiate the quantities in Eq. (14) using different ensemble of observables, e.g., $\phi_{1}(N)$ and $\phi_{2}(N)$.
We tested our hypothesis on SFNs with $\lambda=2.1$, which is the most heterogeneous network considered in our study. In Fig. 3 (a) we use $q=0.7$ to plot $\phi_{1}(N)$ as a function of $N$ for different values of $C$. We could obtain a very clear power-law decay across different values of $q$ and $C$, while consistently observing non-zero values of $\phi_{o}$ when $C=1$. The same explanation holds for results valid for $\phi_{2}$ as well. When we increase the value of $q$, for $C=1$, we find that $\phi_{o,1}$ and $\phi_{o,2}$ decrease since the observables are sampled closer to the transition point and the network becomes sparser. Also, the exponents $\delta_{1}$ and $\delta_{2}$ increase when we increase $q$, regardless of $C$. These findings support that the SPP process with $C>1$ can drastically homogenize the heterogeneous structures in SFNs before the percolation transition happens. Detailed estimations of $\phi_{o,1}$, $\phi_{o,2}$, $\delta_{1}$, and $\delta_{2}$ can be found in Table 2.
Figure 3: Scaling of the $k_{\text{max}}$ ratios. Plot of $\phi_{1}(N)$ as a function of NN with different values of $C$. Results are valid for SFNs with $\lambda=2.1$ using (a) $q=0.7$ (c) $q=0.8$, and (e) $q=0.9$. We also plot $\phi_{k}^{(2)}$ as a function of NN with different values of $C$ using (b) $q=0.7$ (d) $q=0.8$, and (f) $q=0.9$. Plateau values and exponents are reported in Table 2. Symbols are the average values and error bars are the standard deviations. We report the number of realizations used for the scaling analysis in Table S2.
@@ IV.3 Estimation of $1/\bar{\nu}_{1}$ and $1/\bar{\nu}_{2}$
So far, our findings reveal that the ratios of critical exponents $\beta/\bar{\nu}$ and $\gamma/\bar{\nu}$, related to the structure of the network at criticality, depend on whether $C$ is finite or infinite, regardless of the degree exponent $\lambda$. However, we find that the critical exponents $1/\bar{\nu}_{1}$ and $1/\bar{\nu}_{2}$, related to the convergence and fluctuation of pseudo-critical points vary depending on $\lambda$ and $C$ (see Sec. III.3 for details).
@@
First, we report the case when $1/\bar{\nu}=1/\bar{\nu}_{1}=1/\bar{\nu}_{2}$. This is the most common case that has been reported for ordinary percolation on ERNs, random-regular graphs, and low-dimensional lattices with periodic boundary conditions [26, 28]. On SFNs with $\lambda=2.7$, we get $1/\bar{\nu}\approx 0.37$ for finite-$C$ and $1/\bar{\nu}\approx 0.5$ for infinite-$C$, respectively. In figure 4 (b), we plot $|r_{c}-r_{c}(N)|$ and $\sigma_{r_{c}}$ as functions of $N$, respectively, where results are valid for SFNs with $\lambda=2.7$ and SPP with $C=2$ as an example.
@@
Second, when $\lambda>3$, we find that $1/\bar{\nu}_{1}>1/\bar{\nu}_{2}$ for finite-$C>1$, where we get $1/\bar{\nu}_{1}\approx 0.5$ and $1/\bar{\nu}_{2}\approx 0.35$ (See Fig. 4 (c) and (d) for results for $C=2$). Such a relation between $1/\bar{\nu}_{1}$ and $1/\bar{\nu}_{2}$ implies that the convergence of $r_{c}(N)$ towards $r_{c}$ is faster than the fluctuation of $r_{c}(N)$ decreasing as $N$ increases. The explosive percolation is a typical case where this type of discrepancy has been reported [26, 27, 28]. For infinite-$C$, $r_{c}(N)$ initially increases with $N$, but quickly saturates and forms a plateau, remaining nearly constant over a wide range of $N$ (see Fig. S4). Due to this plateau, the quality of least-squares fitting of Eqs. (5) and (12) is compromised. Still, we find that the values of the plateau are within the error bounds of the estimated $r_{c}$. There are two possible scenarios that explain this: One possible explanation is that the size of the network considered is not large enough and the plateau manifests finite-size effects. Another case is when the system size is already large enough; thus, the plateau value represents the true critical point at the infinite-$N$ limit. We confirm that our observation corresponds to the latter, which is supported by a clear data collapse using the obtained $r_{c}$.
@@
Finally, on SFNs with $\lambda=2.1$, we find that $1/\bar{\nu}_{1}<1/\bar{\nu}_{2}$ both for infinite- and finite-$C$. For finite-$C$, we get $1/\bar{\nu}_{1}\approx 0.25$ and $1/\bar{\nu}_{2}\approx 0.44$ while for infinite-$C$, $1/\bar{\nu}_{1}\approx 0.21$ and $1/\bar{\nu}_{2}\approx 0.48$ (see Fig. 4 (a) for results for $C=2$). This is the case reported for bond percolation on high-dimensional lattices above the upper-critical dimension with free-boundary condition [29]. Here, $1/\bar{\nu}_{1}<1/\bar{\nu}_{2}$ implies that the single-instance pseudo-critical points $r_{c}^{*}(N)$ fluctuates within a narrow window, while the pseudo-critical point $r_{c}(N)$ is far away from $r_{c}$ at the infinite-$N$ limit. Results for other values of $C$ can be found in Figs. S5,  S6, and  S7.
@@ IV.4 Data collapse of observables
According to the FSS theory, once we estimate the critical quantities in a fair manner, we could collapse the data across different system sizes by rescaling the data using these critical quantities. To verify our FSS analysis, we display the data-collapse plots in Fig. 5 for SFNs with $\lambda=2.7$ and $k_{\min}=4$, considering the event-based ensemble. By using the estimated critical exponents, we could fairly collapse the data across different orders of magnitude of network size $N$ for the SPP model with different values of $C$. We could also obtain a clear data-collapse considering the conventional ensemble, which supports that our estimation of $r_{c}$ is fair, see Figs. S22,  S23, and  S24. Moreover, the critical exponents imply the same universality class as of ordinary percolation on ERNs, i.e., $\beta/\bar{\nu}=\gamma/\bar{\nu}=1/\bar{\nu}=1/3$.
@@
For SFNs with $\lambda=3.5$ and $\lambda=4.5$, for the SPP model with $C>2$, we find that using $1/\bar{\nu}_{2}$ instead of $1/\bar{\nu}_{1}$ gives better data collapse, regardless of considering the event-based ensemble (Figs. S12,  S13,  S14,  S16,  S17, and  S18) and the conventional ensemble (Figs. S26,  S27,  S28  S30,  S31, and  S32). Especially for finite-$C$, $1/\bar{\nu}_{2}\approx 1/3$ suggests the same universality class of ordinary percolation on ERNs. Note that for infinite $C$, we could achieve a fair data collapse by using $r_{c}$, where the plateau values are within the error bounds (Figs. S28 and  S32). This suggests that the plateau values already converged close enough to the true critical point $r_{c}$. For the case of $C=1$, which corresponds to the ordinary bond percolation, where $1/\bar{\nu}_{1}\approx 1/\bar{\nu}_{2}$, we could obtain a fair data collapse under the event-based ensemble as well as the conventional ensemble, see Figs. S11,  S15,  S25, and  S29.
@@
For SFNs with $\lambda=2.1$, we could obtain a fair data collapse under the event-based ensemble when using $1/\bar{\nu}_{2}$ for finite-$C$ (see Figs. S8 and  S9). For infinite-$C$, the quality of data collapse is compromised (see Fig. S10), but still the data collapse is decent in a narrow range. Under the conventional ensemble, however, we could not obtain a proper data collapse, using both $1/\bar{\nu}_{1}$ and $1/\bar{\nu}_{2}$. In addition, the location where the abscissa value corresponds to zero severely deviates from the peak position of $SN^{-\gamma/\bar{\nu}}$, which corresponds to the pseudo-critical points (see Figs. S19,  S20, and  S21). As suggested from the relation $1/\bar{\nu}_{1}<1/\bar{\nu}_{2}$, we suppose that the pseudo-critical points are still far from the true critical point and a proper FSS behavior is not expected in this regime [29].
@@
Figure 4: Estimation of $1/\bar{\nu}$. We plot $|r_{c}-r_{c}(N)|$ (square symbol) and $\sigma_{r_{c}}$ (circle symbol) as functions of $N$. The slope in the double-log scale corresponds to $1/\bar{\nu}_{1}$ (solid line) and $1/\bar{\nu}_{2}$ (dashed line), respectively. Results are valid for SPP with $C=2$ on SFNs with (a) $\lambda=2.1$, (b) $\lambda=2.7$, (c) $\lambda=3.5$, and (d) $\lambda=4.5$. Symbols are the average values and error bars are the standard deviations. Note that error bars for $\sigma_{r_{c}}$ do not exist so they are omitted in the figure. Estimated critical exponents are reported in Table 1.
@@
Figure 5: Data Collapse of $P$ and $S$. We plot the rescaled observable $PN^{\beta/\bar{\nu}}$ as a function of $r-r_{c}(N)]N^{1/\bar{\nu}}$ for different values of network size $N$. Results are valid for SFNs with $\lambda=2.7$ and $k_{\min}=4$ and SPP model with (a) $C=2$, (b) $C=3$, and (c) $C=\infty$. Similar to (a-c), we plot the rescaled observable $SN^{-\gamma/\bar{\nu}}$ as a function of $(r-\hat{r}_{c})N^{1/\bar{\nu}}$. Note that for SFNs with $\lambda=2.7$, we omit the subscript in $\bar{\nu}_{1}$ and used $1/\bar{\nu}$ since $1/\bar{\nu}_{1}\approx 1/\bar{\nu}_{2}$. The critical exponent ratios used in the figure are reported in Table 1. Note that the range of abscissa is adjusted to display 200 bins of the curve for $N=2^{20}$. We report the number of bins used for each network size in Table S3 for clarity.
@@ V Summary and Discussion
In this work, we systematically investigated the phase transition of SPP on SFNs, considering various degree exponents $\lambda$ and maximum budget $C$. By means of extensive Monte Carlo simulations and FSS analysis, we observed rich behaviors in the SPP transition on SFNs. For $C=1$, which corresponds to the ordinary bond percolation, we obtained results consistent with previous studies [5, 17]. In particular, for $\lambda\leq 3$, the percolation phase transition occurs when all edges are removed from the network; also, for $\lambda\leq 4$, the critical exponents depend on $\lambda$. These anomalous, network-dependent properties are washed out by the SPP process. As long $C>1$, the transition occurs when a critical fraction $r_{c}<1$ of the network edges are removed. For finite $C$ values, we find that the critical exponents describing the structure of the network at criticality are identical to the mean-field exponents on ERNs, i.e., $\beta/\bar{\nu}=\gamma/\bar{\nu}=1/3$, regardless of the $\lambda$. For the infinite-$C$ case, we obtain the critical exponents $\beta/\bar{\nu}\approx 0.2$ and $\gamma/\bar{\nu}\approx 0.6$, which have been observed on ERNs [23, 24], regardless of the value of $\lambda$. These findings lead to the hypothesis that the SPP with $C>1$ drastically homogenizes the heterogeneous structure of SFNs before the SPP transition takes place. We confirmed this hypothesis by analyzing the scaling of the ratio of the initial maximum degree and that before the percolating regime.
@@
The critical exponents $1/\bar{\nu}_{1}$ and $1/\bar{\nu}_{2}$, related to the convergence and fluctuation of the pseudo-critical points, vary depending on $\lambda$ and $C$. To be specific, we find that the relation between $1/\bar{\nu}_{1}$ and $1/\bar{\nu}_{2}$ has a dependency on $\lambda$. When $\lambda=2.1$, we observe $1/\bar{\nu}_{1}<1/\bar{\nu}_{2}$, which has been reported in the case of bond percolation on lattice structure above the upper critical dimension with free boundary condition [29]. However, for $\lambda=2.7$, we observe $1/\bar{\nu}_{1}\approx 1/\bar{\nu}_{2}$, in which most of the ordinary percolation belongs to. Finally, we observe $1/\bar{\nu}_{1}>1/\bar{\nu}_{2}$ for SFNs with $\lambda=3.5$ and $\lambda=4.5$, which is the case of the explosive percolation. Unfortunately, there are only a few studies which report and explain these two exponents [26, 27, 23, 29], thus further investigations on various network structures are needed to better understand such a discrepancy. A recent study also reports rich and non-trivial estimation of these exponents on SFNs depending on the level of the degree cutoff and the degree heterogeneity [30].
@@
Our numerical findings reveal that the SPP process homogenizes the heterogeneous network structure before the percolation transition takes place. However, an analytical description of our numerical findings is still missing. One of the biggest obstacles in formalizing an analytical framework is the topological correlation between edges forming the shortest paths. Unfortunately, typical frameworks such as the one based on generating functions [35, 4] and the one based on message passing [36] assume independence of edges. In future, it would be a promising direction to develop an analytical framework considering “path-based” and “long-ranged” connectivity in percolation.
@@
Finally, we would like to highlight the broad applicability of the SPP framework to various complex systems represented as supply-and-demand networks. These networks include, but are not limited to, transportation networks, communication networks, and supply chain networks, which incorporate resource consumption and depletion. For instance, the SPP has been extended to the minimum-cost percolation and applied to understand the consumption and exhaustion of resources in the U.S. air transportation system while serving the demands [37]. In addition, a similar model to the SPP model has been introduced to study quantum communication networks as well [24]. Since various supply-and-demand network systems are characterized by “scale-free” structure with the existence of hubs, we believe the rich and complex behavior observed in this work can be found in these real-world systems as well.

### v2

@@ IV.2 Shortest-path-percolation process homogenizes the network drastically
where $\left\langle\cdot\right\rangle$ denotes the ensemble average across different realizations, $\phi_{o}$ is the asymptotic limit of the ratio, $\delta$ is the scaling exponent, and $a$ is a constant. Equation (14) implies that if a network still holds its scale-freeness at $r=r^{\dagger}$, we expect $k_{\max}^{\dagger}$ will scale with the size of the network $N$ in the same way as $k_{\max}$ does. In this case, we expect that $\phi_{o}$ is non-zero. On the other hand, we expect to obtain $\phi_{o}=0$ if the SPP process drastically homogenizes the network.
We sample the observables at $r=r^{\dagger}$ using two methods: we collect the observables (i) at a fixed value of $r^{\dagger}=q$ for each realization or (ii) at $r^{\dagger}=qr_{c}^{*}(N)$ for each realization, where $q\in[0,1]$. Here, $q$ is a tunable parameter to decide on how far from the transition point we would like to sample the observables. We choose q=0.7, 0.8,q=0.7,\ 0.8, and $0.9$ to ensure that we are sampling the observables at the supercritical regime. For convenience, we add a subscript to differentiate the quantities in Eq. (14) using different ensembles of observables, e.g., $\phi_{1}(N)$ and $\phi_{2}(N)$.
We test our hypothesis on SFNs with $\lambda=2.1$, which is the most heterogeneous network considered in our study. In Fig. 3 (a) we use $q=0.7$ to plot $\phi_{1}(N)$ as a function of $N$ for different values of $C$. We could obtain a very clear power-law decay across different values of $q$ and $C$, while consistently observing nonzero values of $\phi_{o}$ when $C=1$. The same explanation holds for results valid for $\phi_{2}$ as well. When we increase the value of $q$, for $C=1$, we find that $\phi_{o,1}$ and $\phi_{o,2}$ decrease since the observables are sampled closer to the transition point and the network becomes sparser. Also, the exponents $\delta_{1}$ and $\delta_{2}$ increase when we increase $q$, regardless of $C$. These findings support that the SPP process with $C>1$ can drastically homogenize the heterogeneous structures in SFNs before the percolation transition happens. Detailed estimations of $\phi_{o,1}$, $\phi_{o,2}$, $\delta_{1}$, and $\delta_{2}$ can be found in Table 2. In addition, we perform a complementary analysis and present further evidence showing that the SPP process effectively homogenizes SFNs (see Appendix A for details).
Figure 3: Scaling of the $k_{\text{max}}$ ratios. Plot of $\phi_{1}(N)$ as a function of $N$ for different values of $C$. Results are valid for SFNs with $\lambda=2.1$ using (a) $q=0.7$ (c) $q=0.8$, and (e) $q=0.9$. We also plot $\phi_{2}(N)$ as a function of $N$ for different values of $C$ using (b) $q=0.7$ (d) $q=0.8$, and (f) $q=0.9$. Plateau values and exponents are reported in Table 2. Symbols are the average values and error bars are the standard deviations. We report the number of realizations used for the scaling analysis in Table S2 of Ref. [31].
@@ IV.3 Estimation of $1/\bar{\nu}_{1}$ and $1/\bar{\nu}_{2}$
So far, our findings reveal that the ratios of critical exponents $\beta/\bar{\nu}$ and $\gamma/\bar{\nu}$, related to the structure of the network at criticality, depend on whether $C$ is finite or infinite, regardless of the degree exponent $\lambda$. However, we find that the critical exponents $1/\bar{\nu}_{1}$ and $1/\bar{\nu}_{2}$, related to the convergence and fluctuation of pseudocritical points, vary depending on $\lambda$ and $C$ (see Sec. III.3 for details).
@@
First, we report the case when $1/\bar{\nu}=1/\bar{\nu}_{1}=1/\bar{\nu}_{2}$. This is the most common case that has been reported for ordinary percolation on ERNs, random-regular graphs, and low-dimensional lattices with periodic boundary conditions [26, 28]. On SFNs with $\lambda=2.7$, we get $1/\bar{\nu}\approx 0.37$ for finite $C$ and $1/\bar{\nu}\approx 0.5$ for infinite $C$, respectively. In Fig. 4 (b), we plot $|r_{c}-r_{c}(N)|$ and $\sigma_{r^{*}_{c}}$ as functions of $N$, respectively, where results are valid for SFNs with $\lambda=2.7$ and SPP with $C=2$ as an example.
@@
Second, when $\lambda>3$, we find that $1/\bar{\nu}_{1}>1/\bar{\nu}_{2}$ for finite $C>1$, where we get $1/\bar{\nu}_{1}\approx 0.5$ and $1/\bar{\nu}_{2}\approx 0.35$ [See Figs. 4 (c) and (d) for results for $C=2$]. Such a relation between $1/\bar{\nu}_{1}$ and $1/\bar{\nu}_{2}$ implies that the convergence of $r_{c}(N)$ towards $r_{c}$ is faster than the fluctuation of $r_{c}(N)$ decreasing as $N$ increases. The explosive percolation is a typical case where this type of discrepancy has been reported [26, 27, 28]. For infinite $C$, $r_{c}(N)$ initially increases with $N$, but quickly saturates and forms a plateau, remaining nearly constant over a wide range of $N$ (see Fig. S4 of Ref. [31]). Due to this plateau, the quality of the least-squares fitting of Eqs. (5) and (12) is compromised. Still, we find that the values of the plateau are within the error bounds of the estimated $r_{c}$.Two possible scenarios explain this: One possible explanation is that the size of the network considered is not large enough, and the plateau manifests finite-size effects. Another case is when the system size is already large enough; thus, the plateau value represents the true critical point at the infinite-$N$ limit. We confirm that our observation corresponds to the latter, which is supported by a clear data collapse using the obtained $r_{c}$.
@@
Finally, on SFNs with $\lambda=2.1$, we find that $1/\bar{\nu}_{1}<1/\bar{\nu}_{2}$ both for infinite- and finite-$C$. For finite $C$, we get $1/\bar{\nu}_{1}\approx 0.25$ and $1/\bar{\nu}_{2}\approx 0.44$ while for infinite $C$, $1/\bar{\nu}_{1}\approx 0.21$ and $1/\bar{\nu}_{2}\approx 0.48$ [see Fig. 4 (a) for results for $C=2$]. This is the case reported for bond percolation on high-dimensional lattices above the upper-critical dimension with free-boundary condition [29]. Here, $1/\bar{\nu}_{1}<1/\bar{\nu}_{2}$ implies that the single-instance pseudocritical points $r_{c}^{*}(N)$ fluctuate within a narrow window, while the pseudocritical point $r_{c}(N)$ is far away from $r_{c}$ at the infinite-$N$ limit. Results for other values of $C$ can be found in Figs. S5,  S6, and  S7 of Ref. [31].
@@ IV.4 Data collapse of observables
According to the FSS theory, once we estimate the critical quantities in a fair manner, we can collapse the data across different system sizes by rescaling the data using these critical quantities. To verify our FSS analysis, we display the data-collapse plots in Fig. 5 for SFNs with $\lambda=2.7$ and $k_{\min}=4$, considering the event-based ensemble. By using the estimated critical exponents, we could fairly collapse the data across different orders of magnitude of network size $N$ for the SPP model with different values of $C$. We could also obtain a clear data collapse considering the conventional ensemble, which supports that our estimation of $r_{c}$ is fair, see Figs. S22,  S23, and  S24 of Ref.  [31]. Moreover, the critical exponents imply the same universality class as of ordinary percolation on ERNs, i.e., $\beta/\bar{\nu}=\gamma/\bar{\nu}=1/\bar{\nu}=1/3$.
@@
For SFNs with $\lambda=3.5$ and $\lambda=4.5$, for the SPP model with $C>1$, we find that using $1/\bar{\nu}_{2}$ instead of $1/\bar{\nu}_{1}$ gives better data collapse, regardless of considering the event-based ensemble (see Figs. S12,  S13,  S14,  S16,  S17, and  S18 of Ref. [31]) and the conventional ensemble (Figs. S26,  S27,  S28  S30,  S31, and  S32 of Ref. [31]). Especially for finite $C$, $1/\bar{\nu}_{2}\approx 1/3$ suggests the same universality class of ordinary percolation on ERNs. Note that for infinite $C$, we could achieve a fair data collapse by using $r_{c}$, where the plateau values are within the error bounds (Figs. S28 and  S32 of Ref. [31]). This suggests that the plateau values have already converged close enough to the true critical point $r_{c}$. For the case of $C=1$, which corresponds to the ordinary bond percolation, where $1/\bar{\nu}_{1}\approx 1/\bar{\nu}_{2}$, we could obtain a fair data collapse under the event-based ensemble as well as the conventional ensemble, see Figs. S11,  S15,  S25, and  S29 of Ref. [31].
@@
For SFNs with $\lambda=2.1$, we could obtain a fair data collapse under the event-based ensemble when using $1/\bar{\nu}_{2}$ for finite $C$ (see Figs. S8 and  S9 of Ref. [31]). For infinite $C$, the quality of data collapse is compromised (see Fig. S10 of Ref. [31]), but still the data collapse is decent in a narrow range. Under the conventional ensemble, however, we could not obtain a proper data collapse, using both $1/\bar{\nu}_{1}$ and $1/\bar{\nu}_{2}$. In addition, the location where the abscissa value corresponds to zero severely deviates from the peak position of $SN^{-\gamma/\bar{\nu}}$, which corresponds to the pseudocritical points (see Figs. S19,  S20, and  S21 of Ref. [31]). As suggested from the relation $1/\bar{\nu}_{1}<1/\bar{\nu}_{2}$, we suppose that the pseudocritical points are still far from the true critical point and a proper FSS behavior is not expected in this regime [29].
@@
Figure 4: Estimation of $1/\bar{\nu}$. We plot $|r_{c}-r_{c}(N)|$ (square symbol) and $\sigma_{r^{*}_{c}}$ (circle symbol) as functions of $N$. The slope in the double-log scale corresponds to $1/\bar{\nu}_{1}$ (solid line) and $1/\bar{\nu}_{2}$ (dashed line), respectively. Results are valid for SPP with $C=2$ on SFNs with (a) $\lambda=2.1$, (b) $\lambda=2.7$, (c) $\lambda=3.5$, and (d) $\lambda=4.5$. Symbols are the average values and error bars are the standard deviations. Note that error bars for $\sigma_{r^{*}_{c}}$ do not exist so they are omitted in the figure. Estimated critical exponents are reported in Table 1.
@@
Figure 5: Data collapse of $P$ and $S$. We plot the rescaled observable $PN^{\beta/\bar{\nu}}$ as a function of $r-r_{c}(N)]N^{1/\bar{\nu}}$ for different values of network size $N$. Results are valid for SFNs with $\lambda=2.7$ and $k_{\min}=4$ and SPP model with (a) $C=2$, (c) $C=3$, and (e) $C=N$. Similarly, we plot the rescaled observable $SN^{-\gamma/\bar{\nu}}$ as a function of $r-r_{c}(N)]N^{1/\bar{\nu}}$ for (b) $C=2$, (d) $C=3$, and (f) $C=N$. Note that for SFNs with $\lambda=2.7$, we omit the subscript in $\bar{\nu}_{2}$ and used $1/\bar{\nu}$ since $1/\bar{\nu}_{1}\approx 1/\bar{\nu}_{2}$. The critical exponent ratios used in the figure are reported in Table 1. The range of abscissa is adjusted to display 200 bins of the curve for $N=2^{20}$. We report the number of bins used for each network size in Table S3 of Ref. [31] for clarity.
@@ V Summary and Discussion
In this work, we systematically investigated the phase transition of SPP on SFNs, considering various degree exponents $\lambda$ and maximum budget $C$. By means of extensive Monte Carlo simulations and FSS analysis, we observed rich behaviors in the SPP transition on SFNs. For $C=1$, which corresponds to the ordinary bond percolation, we obtained results consistent with previous studies [5, 17]. In particular, for $\lambda\leq 3$, the percolation phase transition occurs when all edges are removed from the network; also, for $\lambda\leq 4$, the critical exponents depend on $\lambda$. These anomalous, network-dependent properties are washed out by the SPP process. As long as $C>1$, the transition occurs when a critical fraction $r_{c}<1$ of the network edges are removed. For finite $C$ values, we find that the critical exponents describing the structure of the network at criticality are identical to the mean-field exponents on ERNs, i.e., $\beta/\bar{\nu}=\gamma/\bar{\nu}=1/3$, regardless of the value of $\lambda$. For the infinite-$C$ case, we obtain the critical exponents $\beta/\bar{\nu}\approx 0.2$ and $\gamma/\bar{\nu}\approx 0.6$, which have been observed on ERNs [23, 24], regardless of the value of $\lambda$. These findings lead to the hypothesis that the SPP with $C>1$ drastically homogenizes the heterogeneous structure of SFNs before the SPP transition takes place. We confirmed this hypothesis by analyzing the scaling of the ratio of the initial maximum degree and that before the percolating regime.
@@
The critical exponents $1/\bar{\nu}_{1}$ and $1/\bar{\nu}_{2}$, related to the convergence and fluctuation of the pseudocritical points, vary depending on $\lambda$ and $C$. To be specific, for finite $C>1$, we find that the relation between $1/\bar{\nu}_{1}$ and $1/\bar{\nu}_{2}$ has a dependency on $\lambda$. When $\lambda=2.1$, we observe $1/\bar{\nu}_{1}<1/\bar{\nu}_{2}$, which has been reported in the case of bond percolation on lattice structure above the upper critical dimension with free boundary condition [29]. However, for $\lambda=2.7$, we observe $1/\bar{\nu}_{1}\approx 1/\bar{\nu}_{2}$, in which most of the ordinary percolation belongs to. Finally, we observe $1/\bar{\nu}_{1}>1/\bar{\nu}_{2}$ for SFNs with $\lambda=3.5$ and $\lambda=4.5$, which is the case of the explosive percolation. Unfortunately, there are only a few studies which report and explain these two exponents [26, 27, 23, 29]; thus, further investigations on various network structures are needed to better understand such a discrepancy. A recent study also reports rich and non-trivial estimation of these exponents on SFNs depending on the level of the degree cutoff and the degree heterogeneity [30].
@@
Our numerical findings reveal that the SPP process homogenizes the heterogeneous network structure before the percolation transition takes place. However, an analytical description of our numerical findings is still missing. One of the biggest obstacles in formalizing an analytical framework is the topological correlation between edges forming the shortest paths. Unfortunately, typical frameworks such as the one based on generating functions [36, 4] and the one based on message passing [37] assume independence of edges. This limitation also prevents us from fully understanding the boundary of finite and infinite $C$. Possibly, some crossover behavior could occur as $C$ increases beyond some $N$-dependent threshold, which separates the finite-$C$ and infinite-$C$ regimes. In our previous work [23], we numerically verified that the cases $C=N^{1/3}$ and $C=\ln N$ belong to the infinite-$C$ universality class. In the future, it would be a promising direction to develop an analytical framework considering “path-based” and “long-range” connectivity in percolation.
@@
Finally, we would like to highlight the broad applicability of the SPP framework to various complex systems represented as supply-and-demand networks. These networks include, but are not limited to, transportation networks, communication networks, and supply chain networks, which incorporate resource consumption and depletion. For instance, the SPP has been extended to the minimum-cost percolation and applied to understand the consumption and exhaustion of resources in the U.S. air transportation system while serving the demands [38]. In addition, a similar model to the SPP model has been introduced to study quantum communication networks [24]. Since various supply-and-demand network systems are characterized by “scale-free” structure with the existence of hubs, we believe the rich and complex behavior observed in this work can be found in these real-world systems as well.